The ratio of the angle measures of a triangle is 1 : 2 : 3. Find the angle measures. Then classify the triangle by it's angle measures.
Answer by amarjeeth123(575)

The ratio of the angle measures of a triangle is 1 : 2 : 3.
The smallest angle is 1/6*180=30 degrees
The second angle is 2/6*180=60 degrees
The third angle is 3/6*180=90 degrees
The triangle is a right angled triangle.